OpenStudy (anonymous):

the ratio of A to B is a to 8, and the ratio of B to C is 12 to c. if the ratio of A to C is 2 to 1, what is the ratio of a to c?

OpenStudy (anonymous):

Are A and C different from a and c? Maybe angles and sides?

OpenStudy (anonymous):

theyre different yes

OpenStudy (anonymous):

Is there a diagram that goes along with the problem?

OpenStudy (anonymous):

no diagram

OpenStudy (anonymous):

i think they stand for angles and sides like you said before

OpenStudy (anonymous):

OH, I misread. I missed that "12 to c"

myininaya (myininaya):

A/B=a/8 B/C=12/c A/c=2/1 what is a/c? A/B * 8=a B/C *c=12 => c=12*C/B a/c=(A/B * 8)/(12* C/B)=8/12 * A/C=2/3 * A/C

myininaya (myininaya):

but A/C=2/1=2 so a/c=2/3 * 2=4/3